# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v2

@@ 5.1 Experiment Settings
For link prediction on HKGs, we select three benchmark datasets: WikiPeople Guan et al. (2021) , WD50K Galkin et al. (2020) and JF17K Wen et al. (2016). As for the NKGs, we choose FBH, FBHE and DBHE constructed by Chung and Whang (2023). Lastly, we use wikidata12k Dasgupta et al. (2018), a subset of wikidata Vrandečić and Krötzsch (2014) for link prediction on TKGs. The statics of datasets are given in Appendix D.
@@
We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts. We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. And we choose to adopt the metrics used in prior works. Additionally, we employ filtering settings Bordes et al. (2013) during the evaluation process to eliminate existing facts in the dataset.
@@
For link prediction on HKG, we compare our UniHR against NaLP Guan et al. (2021), tNaLP Guan et al. (2021), RAM Liu et al. (2021), HINGE Rosso et al. (2020), NeuInfer Guan et al. (2020), StarE Galkin et al. (2020), HyTransformer Yu and Yang (2021), GRAN Wang et al. (2021) and HyNT Chung et al. (2023). For link prediction on NKG, QuatE Zhang et al. (2019), BiQUE Guo and Kok (2021), Neural-LP Yang et al. (2017), DRUM Sadeghian et al. (2019), AnyBURL Meilicke et al. (2019), BiVE Chung and Whang (2023) and NestE Xiong et al. (2024) are chosen as baselines. BiVE and NestE are especially designed for NKG. We compare against following TKG link prediction methods: ComplEx-N3 Lacroix et al. (2018), TTransE Leblay and Chekol (2018), HyTE Dasgupta et al. (2018), TA-DistMult Garcia-Duran et al. (2018), ATiSE Xu et al. (2019), TeRo Xu et al. (2020), TASTER Wang et al. (2023), TGeomE+ Xu et al. (2023a).
@@
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from Chung and Whang (2023) for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}0$ to prevent overfitting. We employ AdamW Kingma and Ba (2015) optimizer. Hyperparameters are chosen by using a grid search based on the MRR performance and details can be found in Appendix F.
@@ 5.2 Main Results
Table 3: Results of link prediction on wikidata12k.
@@
# Tr. Params MRR Hits@1 Hits@3 Hits@10
ComplEx-N3 - 0.248 0.143 - 0.489
TTransE - 0.172 0.096 0.184 0.329
HyTE 1.69M 0.253 0.147 - 0.483
TA-DistMult - 0.230 0.130 - 0.461
TeRo 12.91M 0.299 0.198 0.329 0.507
ATiSE 31.46M 0.252 0.148 0.288 0.462
TASTER 5.04M 0.327 0.239 - 0.514
TGeomE+ - 0.333 0.232 0.361 0.546
UniHR 3.68M 0.333 0.240 0.367 0.527
Table 4: Ablation studies on the HiSL module for triple prediction. Best results are boldfaced.
@@
Model WikiPeople- wikidata12k-
subject/object all entities subject/object
MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10
@@
Link prediction on HKG.
We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 1. We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT) demonstrate significantly better performance compared to other models. We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts. Among these methods, it can be seen that our proposed UniHR achieves state-of-the-art results, which means our method effectively captures global structural information. Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K. This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information. Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts, so our training parameters do not significantly increase.
Link prediction on NKG.
Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction. Unlike previous quaternion-based methods Chung and Whang (2023); Xiong et al. (2024), UniHR significantly reduces the number of required training parameters. From the results in Table 2, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs. For base link prediction task, UniHR achieves considerable improvements on all datasets. Of particular note, the MRR of FBHE/FBH and DBHE increases by 8.1% and 3.8%, respectively.
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do. We will further illustrate the effectiveness of UniHR with other decoders in Appendix C.
Link prediction on TKG.
As shown in Table 3, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3. However, existing temporal knowledge graph embedding methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs. In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
5.3 Further Analysis
Ablation study on HiSL.
We conduct ablation experiments on triple prediction, the most relevant task to fact nodes, using three NKG datasets. As shown in Table 4, all variant models with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$) in HiSL. There is a performance decrease, indicating that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
Potential of Joint Learning on Different Types of KGs.
We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs. To explore its potential, we conduct joint learning on different types of KGs. Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata Vrandečić and Krötzsch (2014), namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations. Due to the different types of facts, there are no identical facts in these two subsets. To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set. Statics of wikimix are given in Appendix D.
@@ 5.3 Further Analysis
6 Conclusion
@@
In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Our approach achieves the best or competitive performance on link prediction tasks across three types of KGs.
Limitations
The limitations of our paper are summarized as follows:
Our UniHR
In this paper, our UniHR framework focuses on link prediction tasks under transductive settings with a single modality. In the future, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning Teru et al. (2020) and multi-modal scenarios Zhang et al. (2024), etc.
Joint Learning on Different Types of KGs
@@
Constrained by computational resources, our analysis of the potential for joint training across multiple types of knowledge graphs focuse only on HKG and TKG. We believe the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs. In the future, we aim to explore joint training across more types of KG to demonstrate the advantages of integrating multi-type KG data.
Ethics Statement
In this paper, we explore the unified knowledge graph link prediction problem, aiming to complete various types of knowledge graphs using a unified model with deep learning techniques. Our training and evaluation are based on publicly available and widely used datasets of different types of knowledge graphs. Therefore, we believe this does not violate any ethics.
Appendix A Related Works
A.1 Link Prediction Methods for Triple-based KGs

### v3

@@ 5.1 Experiment Settings
For link prediction on HKGs, we select three benchmark datasets: WikiPeople [15] and WD50K [12]. For link prediction on NKGs, we choose FBH, FBHE and DBHE constructed by [7]. For link prediction on TKGs, we use wikidata12k [8], a subset of wikidata [33]. To further evaluate the potential of the unified representation, we further introduce the hyper-relational TKG datasets WIKI-hy and YAGO-hy [10]. The detailed statics are given in Table 1.
Table 1: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)" column denotes the number of atomic facts in nested facts.
Dataset Fact Entities Rela Train Valid Test with Q(%) Arity N-Fact N-Rela AF(%) Period
Hyper-relational Knowledge Graph
WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
Nested Factual Knowledge Graph
FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
Temporal Knowledge Graph
wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
Hyper-relational Temporal Knowledge Graph
WiKi-hy 139078 16634 147 111252 13900 13926 13335(9.59%) 2-8 - - - 1513-2020
YAGO-hy 73143 16167 54 51193 10973 10977 5107(6.98%) 2-5 - - - 0-187
Multiple types of Knowledge Graph
wikimix 409566 43832 184 326936 41777 3525/37328 9098(2.2%) 2-7 - - - 19-2020
@@
We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts. We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings [4] during the evaluation to eliminate existing facts in the dataset. It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set. Similarly, the difference in filtering settings of TKG occurs in timestamp. In this paper, we adopt the strict filtering setting of the former. To ensure fair comparison, for HKG we utilize the results of HyNT [6] with the same settings as ours. For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
@@
For link prediction on HKG, we compare our UniHR with NaLP [15], tNaLP [15], RAM [22], HINGE [27], NeuInfer [14], StarE [12], HyTransformer [43], GRAN [34], HyNT [6] and HAHE [23]. For link prediction on NKG, QuatE [45], BiQUE [16], Neural-LP [42], DRUM [28], AnyBURL [24], BiVE [7] and NestE [37] are chosen as baselines. Among them, BiVE and NestE are especially designed for NKG. We compare against following TKG link prediction methods: ComplEx-N3 [20], HyTE [8], TA-DistMult [13], ATiSE [38], TeRo [39], TGeomE+ [40], HGE [26] and ECEformer [11].
@@
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from [7] for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting. We employ AdamW [19] optimizer. Hyperparameter details can be found in Appendix C.
@@ 5.2 Main Results
Table 2: Results of link prediction on HKG datasets. Results are reported by us, and the others are taken from [6]. Best results are bold, while the second are underlined.
WikiPeople WD50K
Model subject/object all entities subject/object all entities
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
NaLP [15] 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
tNaLP [15] 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
RAM [22] 0.459 0.384 0.584 0.461 0.386 0.585 0.276 0.210 0.399 0.296 0.232 0.416
HINGE [27] 0.393 0.309 0.547 0.395 0.311 0.549 0.264 0.187 0.410 0.277 0.200 0.424
NeuInfer [14] 0.357 0.247 0.533 0.357 0.248 0.532 0.220 0.154 0.347 0.225 0.158 0.355
StarE [12] 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
HyTransformer [43] 0.465 0.395 0.598 - - - 0.308 0.239 0.452 - - -
GRAN [34] 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
HyNT [6] 0.482 0.415 0.602 0.481 0.414 0.603 0.333 0.259 0.474 0.360 0.287 0.500
HAHE [23] 0.498 0.426 0.610 0.497 0.429 0.614 0.343 0.270 0.484 0.378 0.306 0.515
UniHR 0.491 0.417 0.618 0.493 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.513
Link Prediction on HKG.
We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 2. We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT, HAHE) demonstrate significantly better performance compared to other models. We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts. Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE, which means our method effectively captures hierarchical fact information. Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K. This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information. Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts through simple MLPs, so our training parameters do not significantly increase.
Table 3: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from [37]. We pick the best variants of BiVE [7] and NestE [37].
FBH DBHE FBH FBHE DBHE
Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
Base link prediction Triple prediction
QuatE [45] 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 0.315 0.486 0.233 0.357 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [12] 0.317 0.490 0.237 0.359 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 0.310 0.526 0.220 0.364 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
UniHR 0.401 0.619 0.296 0.448 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
Link Prediction on NKG.
Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction. From the results in Table 3, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs. For base link prediction task on triple-based KGs, UniHR achieves considerable improvements. Of particular note, the MRR of FBHE increases by 8.1%.
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do. We will further illustrate the effectiveness of UniHR equipped with other decoders in Appendix B.
Table 4: Results of link prediction on wikidata12k. Results are reported by us, and others are taken from original papers.
@@
MRR H@1 H@3 H@10
ComplEx-N3 [20] 0.248 0.143 - 0.489
HyTE [8] 0.253 0.147 - 0.483
TA-DistMult [13] 0.230 0.130 - 0.461
TeRo [39] 0.299 0.198 0.329 0.507
ATiSE [38] 0.252 0.148 0.288 0.462
TGeomE+ [40] 0.333 0.232 0.361 0.546
HGE [26] 0.290 0.176 0.323 0.514
ECEformer [11] 0.262 0.159 0.255 0.462
UniHR 0.333 0.240 0.367 0.527
Link Prediction on TKG.
As shown in Table 4, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3. However, existing temporal knowledge graph embedding (KGE) methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs. In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
Table 5: Ablation studies on triple prediction task.
@@
5.3 Ablation Study on HiSL
We conduct ablation experiments on triple prediction, the most relevant task to fact nodes. As shown in Table 5, all variants with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
5.4 Potential of Unified Representation
Table 6: Results on hyper-relational TKG datasets.
Model WiKi-hy YAGO-hy
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
TeRo [39] 0.572 0.473 0.640 0.727 0.760 0.720 0.782 0.822
HGE [26] 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
StarE [12] 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
GRAN [34] 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
HyNT [6] 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
HypeTKG [10] 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
UniHR 0.681 0.618 0.717 0.774 0.823 0.801 0.839 0.859
Generalize to Hyper-relational TKGs.
Due to the unity of representation, UniHR can easily generalize to compositional types of KGs, like hyper-relational TKGs (HTKGs) [10] that integrate HKGs and TKGs. Every H-fact in the HTKG is coupled with a timestamp explicitly indicating its time validity. From results in Table 6, UniHR offers a significant performance improvement in link prediction tasks on hyper-relational TKGs compared to TKG-specific and HKG-specific models, demonstrating a good collaborative modeling ability for auxiliary key-value pairs and timestamps. Our UniHR even achieves competitive performance with the specialized hyper-relational TKG model HypeTKG without complex module stacking.
Table 7: Results of separate and joint training on NKG.
FBH DBHE FBH DBHE
Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
Base link prediction Triple prediction
UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
UniHRJoint 0.402 0.632 0.298 0.452 0.948 0.993 0.857 0.989
Joint Learning on Different Tasks of KGs.
For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form. Therefore, we attempt joint training on the two tasks using the NKG dataset, as shown in Table 7. We observe that the results of joint training are generally superior to those of separate training, indicating that nested facts and atomic facts can mutually enhance and supplement the semantics.
@@ 5.4 Potential of Unified Representation
WikiPeople- wikidata12k-
Model subject/object all entities subject/object
MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
@@
Joint Learning on Different Types of KGs.
We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs. To explore its potential, we conduct joint learning on different types of KGs. Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata [33], namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations. Due to the different types of facts, there are no identical facts in these two subsets. To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set. Statics of wikimix are given in Table 1.
@@
Table 9: Number of training parameters and times.
Dataset Model Training Params Training Times
WD50K StarE [12] 10.81M 48h+
HyNT [6] 29.61M 22.3h
HAHE [23] 42.35M 29.2h
UniHR 10.55M 23.4h
wikidata12k ATiSE [38] 31.46M 5.8h
HGE [26] 3.44M -
ECEformer [11] 11.99M 5.6h
UniHR 3.68M 3.6h
YAGO-hy HypeTKG [10] 10.83M 37.5h
UniHR 4.37M 17.6h
5.5 Efficiency Analysis
Compared to existing state-of-the-art methods, UniHR does not significantly increase the overhead in terms of model training parameters and runtime as shown in Figure 9. The embeddings of relation nodes and fact nodes are computed using a simple MLP based on atomic nodes and relations, which does not significantly increase the number of training parameters. For runtime, we do not use the whole graph during all message passing phases and adopt a dropout strategy to prevent overfitting, which can effectively enhance training efficiency. Therefore, our UniHR achieves a superior trade-off between model effectiveness and efficiency.
6 Conclusion and Future Work
@@
In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Our UniHR achieves the best or competitive performance across five types of KGs, and highlight the strong potential of unified representations.
@@
In the future, we believe that the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs. Moreover, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning [30] and multi-modal scenarios, etc.
Supplemental Material Statement:
The dataset, code, and hyperparameter details are available at https://github.com/Lza12a/UniHR.
Appendix
Appendix 0.A Related Works
0.A.1 Link Prediction on Triple-based KGs